# Uncovering Bias in Large Vision-Language Models at Scale with Counterfactuals

### v1

@@ Appendix I Numeric Question Analysis
tattooed White []
young Asian studious
young Black []
young Indian []
@@
young Middle Eastern []
@@
Table 25: Analysis of stereotypes in GPT-4o generations on physical-race subset (“Characteristics” prompt only).
@@
Black male []
Black female []
Asian male studious, martial, artist, scientist
Asian female []
Middle Eastern male []
Middle Eastern female hospitable, modest
Latino male manual, labor, serving, blue-collar
Latino female []
Indian male dancer, shopkeeper, security, policeman, guard, clerical
Indian female []
@@
White female blonde, blond, caregiver
Table 26: Analysis of stereotypes in GPT-4o generations on the race-gender subset (“Characteristics” prompt only).
@@
Label Prompt
Table 29: Prompts used to generate numerical responses.

### v2

@@ Appendix H Inference-Time Mitigation
tattooed White rebellious, individualistic, rebellion, alternative, rebel, motorcyclist, intimidating, edgy, oppressive, tribal, graffiti, augmented, gangs, underground, gritty, blue-collar, troubled, corrupt, tough, rough, gang, unethical, rogue, conventional, corruption, hacking
young Asian calligraphy, tea, samurai, kendo, martial
young Black afro, basketball, hip-hop, chores, jazz
young Indian cricket, schoolwork, turban, homework, rural, low-cost, farmer, middle-class, grocery, agricultural, homemaker, secondary, yoga, headscarf, tutoring, entrepreneur, educational, over-the-ear, teachers, solar-powered, boy, affordable, chess, extracurricular, villages
@@
young Middle Eastern headscarf, underdeveloped
@@
Table 25: Analysis of stereotypes in GPT-4o generations on physical-race subset.
@@
Black male basketball, hip-hop, saxophone, football, jazz, athlete, muscular, trouble, mechanic, working-class, homeless, poverty, sports, gang-related, laborer
Black female braids, afro, curly, nigeria, african, kenya, harlem, caribbean, malaria, africa, creole, cuban, louisiana, jazz
Asian male sensei, dragon, dojo, warlord, samurai, karate, martial, zen, monks, chi, grandmaster, lotus, wushu, judo, clan, calligraphy, philosopher, aikido, philosophies, taekwondo, monastery, warriors, meditative, tournaments, ceremonies, agility, agile, exercises, chinese, defend, disciplines, contemplative, jungle, masters, fortitude, meditating, speed, asian, philosophical, opponent, demonstrations, crane, warrior, combat, mastery, fighting, arts, artist, vigilante, tea, immigrants, capoeira, programmers
Asian female teahouse, calligrapher, calligraphy, hairpiece, pottery, wushu, tea, flower, cherry, ceremony, embroidery, lotus, ceremonies, dragon, costume, blossom, kendo
Middle Eastern male bearded, tribe, beard, desert, desert-like, arid, dunes, arabic, dark-haired, palm, turban, sands, bandits, merchants, traders, shirtless, warriors, warrior, eastern, veil, east, marketplace, spice
Middle Eastern female harem, desert, mystical, modesty, conservative
Latino male muscular, tattoo, well-toned, mexican, argentine, soccer, spanish, brazilian, football, flamenco
Latino female flamenco, puerto, tango, rico, latin, fiery, mexican, bodega, cuban, havana, argentina
Indian male dhoti, mustache, cricket, sitar, kathak, homemaker, turban, beret, bazaar, bracelets, irrigation, bollywood, guru, tabla, ganges, anklet, sanitation, spirituality, shopkeeper, farmer, chai, divine, villagers, sweets, agricultural, enlightenment, ayurvedic, village, gods, bharatanatyam, temple, temples, spice, rural, anklets, traditionalists, garments, mridangam, carnatic, yoga, serpent, festivals, jewelry, classical, spices, corruption, bureaucratic, craftsmen, farmers, poverty, ornate
Indian female sari, saree, bangles, armlets, sarees, bangle, jewelry, necklaces, carnatic, blouse, saris, mudras, bharatanatyam, bindi, anklets, dhoti-style, natya, lehenga, ghungroos, sari-like, ghungroo, gold, braid, bracelet, goddess, temples, bracelets, chai, embroidery, gender, pendant, flowers, waistband, patterns, ornament, conservative, adorns, costume, festivals, adorn, adorned, tabla, outfit, costumes, sitar, decorative, traditionalists, jasmine, cosmetics, embroidered, colored, beautiful, singer
@@
White female blonde, blond, leotard, ballerinas, ballerina, tutu, arabesques, prima, madame, arabesque, twirling, soloist, slippers, nutcracker, pirouettes, pirouette, ballet, technology-savvy, girl, hygienic
Table 26: Analysis of stereotypes in GPT-4o generations on the race-gender subset.